# Graph Condensation: A Survey

### v1

@@ 2.3 Taxonomy
The above criteria not only serve as the indicators of GC’s utility, but also act as the objectives that steer extensive research efforts in this field. To streamline the understanding of GC, we class GC methods into four categories which align with the above criteria: effective GC, generalized GC, fair GC, and efficient GC, as shown in Figure 1.
In addition to our proposed taxonomy, we also provide a detailed summary of GC methods, highlighting their distinct characteristics, in Table 1. This table offers a comprehensive comparison across several dimensions, including motivation, optimization strategy, condensed graph generation method, and downstream tasks.
Table 1: Summary of existing graph condensation research, organized chronologically by publication date. In optimization strategy: “GM” represents gradient matching, “DM” refers to distribution matching, “TM” is short for trajectory matching and “KRR” is kernel ridge regression. In relay model: “GNN” is graph neural network, “GNTK” refers to graph neural tangent kernel and “SD” is short for spectral decomposition. In downstream task: “NC” represents node classification, “GC” refers to graph classification, “AD” is anomaly detection and “LP” is short for link prediction.
Method Motivation Input Optimization strategy Output Downstream task Enhanced module
Pre-processing Objective Relay model Bi-level Structure construction Sparsification
GCond Jin et al. (2022b) Effectiveness GM GNN Generative model Threshold NC N/A
DosCond Jin et al. (2022a) Efficiency GM GNN Parameterization Gumbel softmax GC Optimization
GCDM Liu et al. (2022) Efficiency DM GNN Generative model Threshold NC Objective
HCDC Ding et al. (2022) Application GM GNN Generative model Threshold NC Condensed graph
MSGC Gao and Wu (2023) Effectiveness GM GNN Pre-defined structure N/A NC Condensed graph
SGDD Yang et al. (2023a) Generalization GM GNN Generative model Regularization NC, AD, LP Objective
SFGC Zheng et al. (2023c) Effectiveness TM GNN Pre-defined structure N/A NC Objective
FGD Feng et al. (2023) Fairness GM GNN Generative model Threshold NC Objective
GCARe Mao et al. (2023) Fairness GM GNN Generative model Threshold NC Relay model
MCond Gao et al. (2024) Application GM GNN Generative model Threshold NC Objective
KIDD Xu et al. (2023) Efficiency KRR GNTK Parameterization Gumbel softmax GC Relay model
CaT Liu et al. (2023b) Application DM GNN Pre-defined structure N/A NC Optimization
CTRL Anonymous (2023) Effectiveness GM GNN Generative model Threshold NC, GC Objective
FedGKD Pan et al. (2023) Application N/A GNN Generative model Gumbel softmax NC Condensed graph
GCEM Liu et al. (2023a) Generalization DM SD Parameterization N/A NC Relay model
GC-SNTK Wang et al. (2023) Efficiency KRR GNTK Parameterization N/A NC Relay model
GroC Li et al. (2023) Effectiveness GM GNN Generative model Threshold NC, GC Objective
PUMA Liu et al. (2023c) Application DM GNN Pre-defined structure N/A NC Objective
@@ 2.3.1 Effective Graph Condensation
The primary objective of GC focuses on optimizing the task-specific quality of the condensed graph. As illustrated in Figure 3, the condensed graph produced by effective GC methods enables GNNs to achieve performance that closely approximates those trained on the original graph, especially in targeted downstream tasks. Consequently, advanced effective GC significantly enhances the feasibility of using condensed graphs as substitutes. To systematically study these methods, we further divide this category into two distinct classes: optimization-driven methods and module-enhanced methods, providing a clearer and more detailed exploration of effective GC methodology.
@@ 2.3.2 Generalized Graph Condensation
The objective of GC extends beyond merely training specific models for designated tasks. It also involves enabling the application of various models across a spectrum of tasks, as depicted in Figure 4. A condensed graph with robust generalization capabilities avoids repeated generation of task-specific or model-specific condensed graphs, thereby significantly enhancing its utility in various applications, such as neural architecture search, hyper-parameter optimization, and multi-task learning. The primary challenge of generalization is to minimize the loss of information in both the structural and feature representations during condensation. Consequently, we study existing generalized GC from two primary aspects: advanced relay model for better information preservation and regularization technique to compensate information loss during condensation.
@@ 2.3.3 Fair Graph Condensation
Balancing information during the condensation process represents a pivotal challenge in the domain of GC. The intensive compression of GC often leads to an amplification of data biases within the condensed graphs. Consequently, GNNs trained on such condensed graphs exhibiting more pronounced fairness issues compared to those trained on original graphs. In response to this, fair GC methods constrain the condensation procedure to mitigate bias in the condensed graph as shown in Figure 5. The approaches to improving fairness in GC encompass two key ideas. The first idea is to integrate the fairness-enhancing regularization into the GC objective function, effectively guiding the optimization process toward fairer outcomes. The second idea is the development of novel modules within the relay model, promoting unbiased representations in the condensed graphs.
@@
2.3.4 Efficient Graph Condensation
An efficient condensing procedure is essential for rapid adaptation in life-long graph learning scenarios. However, GC often suffers from sophisticated optimization process and slow convergence, resulting in a time-consuming condensing procedure. This particularly deteriorates in life-long learning contexts, where regular updates to the condensed graph dataset are a common requirement, as depicted in Figure 6. To address these challenges, efficient GC methods have been developed to accelerate the condensation process across all components within GC . To thoroughly explore the mechanisms underlying efficient GC, we break down the conventional GC procedure into three stages: graph encoding, optimization, and graph generation. These components align with the input, processing, and output stages of GC, respectively. By enhancing the efficiency of each stage, efficient GC methods significantly reduce the time required for the entire GC process.
@@ 3 Optimization Strategies
As we delve deeper into the intricacies of GC, it is crucial to examine the optimization strategies that form a core module within the GC framework. We hereby introduce the various optimization strategies employed in GC, offering a high-level perspective on the condensation procedure. Our discussion commences with the foundational principle of optimization objective (Eq. (3)), then move onto detailed optimization strategies (Eq. (4)-(8)).
The foundational principle of the optimization strategies posits that the performance of a model trained on the condensed data should align with that of a model trained on the original data. To this end, we initially define the loss for data $\mathcal{T}$ and $\mathcal{S}$ in relation to the parameter ${\theta}$ of the relay model as:
@@
$\min_{\mathcal{S}}\mathcal{L}^{\mathcal{T}}\left(\theta^{\mathcal{S}}\right)\>\>\text{s.t.}\>\>\theta^{\mathcal{S}}=\arg\min_{\mathcal{\theta}}\mathcal{L}^{\mathcal{S}}\left(\theta\right),$ (3)
where the inner loop optimizes the parameter $\theta$ of relay model on the condensed data until convergence, and the outer loop subsequently optimizes the condensed data $\mathcal{S}$. Addressing the above objective necessitates the incorporation of the task-specific objective $\ell$ and entails resolving a nested loop optimization, which includes unrolling the entire training trajectory of the relay model optimization. This process can be prohibitively expensive. To alleviate the complexity of the optimization process, various approximation methods have been proposed, including gradient matching, trajectory matching, kernel ridge regression and distribution matching.
Gradient matching. Gradient matching is initially proposed by Zhao et al. (2020) and formulates the condensation objective in Eq. (3) as matching the optimized parameters of the models trained on two datasets:
@@
where $\theta_{0}$ is the initialization of ${\theta}^{\mathcal{T}}$ and ${\theta}^{\mathcal{S}}$. ${\Theta}$ is a specific distribution for relay model initialization. The expectation on $\theta_{0}$ aims to improve the robustness of ${\mathcal{S}}$ to different parameter initialization Lei and Tao (2024). $\mathcal{D}\left(\cdot,\cdot\right)$ is a function measuring the distance between the gradients w.r.t. ${\theta}^{\mathcal{T}}$ and ${\theta}^{\mathcal{S}}$. The bi-level objective in Eq. (3) is approximated by matching the model gradients at each training step $t$. In this way, the training trajectory on condensed data can mimic that on the original data, i.e., the models trained on these two datasets converge to the similar solutions. The optimization objective $\mathcal{L}_{cond}$ in Eq. (1) is defined as:
$\begin{split}\mathcal{L}_{cond}=\mathbb{E}_{\theta_{0}\sim{\Theta}}&\left[\sum_{t=1}^{T}\mathcal{D}\left(\nabla_{\theta}\mathcal{L}^{\mathcal{T}}\left({\theta_{t}}\right),\nabla_{\theta}\mathcal{L}^{\mathcal{S}}\left({\theta_{t}}\right)\right)\right]\,\,\\ &\text{s.t.}\,\,{\theta}_{t+1}=\operatorname{opt}\left(\mathcal{L}^{\mathcal{S}}\left({\theta_{t}}\right)\right),\end{split}$ (5)
@@
Trajectory matching. Gradient matching primarily aligns single-step gradients, yet it may accumulate errors when the relay model is iteratively updated using condensed data across multiple steps. To mitigate this problem, Cazenavette et al. (2022) propose a multi-step matching approach to approximate Eq. (4) and known as trajectory matching.
Trajectory matching trains two separate relay models on condensed data and original data, respectively. These two relay models are trained for distinct steps starting from a same initialization $\theta^{\prime}_{t}$, which is sampled from the original data training checkpoints ${\Theta^{\prime}}$. Then, trajectory matching aims to minimize the discrepancy between the final points of these two training trajectories:
@@
Kernel ridge regression. Gradient matching and trajectory matching approximately optimize the objective Eq. (4) by matching one-step gradient or multi-step trajectory. However, these inexact solutions can hardly lead to the optimal condensed graph. To bridge this gap, Nguyen et al. (2020) transform the original task into a regression problem by substituting the neural network with kernel ridge regression (KRR). This approach enables the direct resolution of the constraint in Eq. (3) using a closed-form solution. Consequently, the complex nested optimization is simplified into a first-order process by considering the condensed data as the support set and the original data as the target set. The optimization problem described in Eq. (3) is reformulated into a regression task:
@@
Distribution matching. All aforementioned optimization strategies stem from the objective of aligning model performance as outlined in Eq. (3), inevitably coupling the relay model with the optimization process. To avoid this issue, Zhao and Bilen (2023) propose an alternative optimization approach known as distribution matching, which focuses on matching the feature distributions of condensed and original data. The objective of distribution matching is to generate condensed data with a feature distribution closely approximating that of the original data. Accordingly, the optimization objective of distribution matching is formulated as:
@@
To address the inconsistency in sample quantities between original and condensed dataset, the distance function $\mathcal{D}\left(\cdot,\cdot\right)$ is specifically designed for class-wise comparison. Moreover, distribution matching eliminates the requirement for the task-specific objective in Eq. (2), thereby enhancing the efficiency of the condensation process.

### v2

@@ 2.3 Taxonomy
The above criteria not only serve as indicators of utility for GC but also act as the objectives that steer the extensive research efforts in this field. To streamline the understanding of GC, we class GC methods into five categories that align with the above criteria: effective GC, generalized GC, efficient GC, fair GC, and robust GC, as shown in Figure 3. The paper distribution for each category is show in Figure LABEL:figpie (a).
@@ 2.3.1 Effective Graph Condensation
The primary objective of GC focuses on optimizing the task-specific quality of the condensed graph. As illustrated in Figure 5, the condensed graph produced by effective GC methods enables GNNs to achieve performance that closely approximates those trained on the original graph, especially in targeted downstream tasks. To systematically study these methods, we further divide this category into two distinct classes: optimization-driven methods and augmentation-based methods, facilitating an in-depth understanding of techniques in effective GC.
@@ 2.3.2 Generalized Graph Condensation
The objective of GC extends beyond merely training specific models for designated tasks. It also involves enabling the application of various models across a spectrum of tasks, as depicted in Figure 6. A condensed graph with good generalization capabilities avoids the repeated generation of task-specific or model-specific condensed graphs, thereby significantly enhancing its utility in various applications, such as open-world graph learning [70], neural architecture search [12] and hyper-parameter optimization [71]. The primary challenge of generalization is to capture the key information in both the structural and feature representations during condensation. Consequently, we study existing generalized GC from two primary aspects: the spectral and spatial techniques for better information preservation during condensation.
@@
2.3.3 Efficient Graph Condensation
An efficient condensing procedure is essential for rapid adaptation in life-long graph learning scenarios. However, GC often suffers from sophisticated optimization processes and slow convergence, resulting in a time-consuming procedure. This particularly deteriorates in life-long learning contexts, where regular updates to the condensed graph dataset are a common requirement, as depicted in Figure 7. To address these challenges, efficient GC methods have been developed to accelerate the condensation process across all components within GC. To thoroughly explore the mechanisms underlying efficient GC, we break down the conventional GC procedure into three stages: graph encoding, optimization, and graph generation. These components align with the input, processing, and output stages of GC, respectively. By enhancing the efficiency of each stage, efficient GC methods significantly reduce the time required for the entire GC process.
@@ 2.3.4 Fair Graph Condensation
Balancing information during the condensation process represents a pivotal challenge in the domain of GC. The intensive compression of GC often leads to an amplification of data biases within the condensed graphs. Consequently, GNNs trained on such condensed graphs exhibit more pronounced fairness issues compared to those trained on original graphs. In response to this, fair GC methods constrain the condensation procedure to mitigate bias in the condensed graph as shown in Figure 8. The approaches to improving fairness in GC encompass two key aspects: the first integrates the fairness-enhancing regularization into the GC objective function, effectively guiding the optimization process toward fairer outcomes. The second aspect is the development of advanced relay model, promoting unbiased representations in the condensed graphs.
@@
2.3.5 Robust Graph Condensation
A robust condensation procedure is significant for unskewed condensed graph generation and deployment in noisy real-world scenarios. Conventional GC indiscriminately emulates the original graph distributions by condensed graph, disregarding the intrinsic quality of the graph. As a result, the condensed graph inherits the noisy distribution from the original graph, which is carried over to GNNs and compromises the model prediction accuracy. To address this issue, Robust GC aims to filter out the noise within the original graph during the condensation procedure and extracts only the core, causal information to the condensed graph for effective GNN training, as depicted in Figure 9.
Figure 9: Robust graph condensation (GC) ensures that GNNs trained on condensed graphs produce comparable results relative to those from GNNs trained on the clean original graph.
@@ 3 Optimization Strategies
As we delve deeper into the intricacies of GC, it is crucial to examine the optimization strategies that form a core module within the GC framework. We hereby introduce the various optimization strategies employed in GC, offering a high-level perspective on the condensation procedure. Our discussion commences with the foundational principle of optimization objective (Eq. (4)), then moves onto detailed optimization strategies (Eq. (5)-(9)).
The foundational principle of the optimization strategies posits that the performance of a model trained on the condensed data should align with that of a model trained on the original data. To this end, the losses for data $\mathcal{T}$ and $\mathcal{S}$ in relation to the parameter ${\theta}$ of the relay model are initially defined as:
@@
$\min_{\mathcal{S}}\mathcal{L}^{\mathcal{T}}\left(\theta^{\mathcal{S}}\right)\>\>\text{s.t.}\>\>\theta^{\mathcal{S}}=\arg\min_{\mathcal{\theta}}\mathcal{L}^{\mathcal{S}}\left(\theta\right).$ (4)
Addressing the objective outlined in Eq. (4) necessitates the incorporation of the task-specific objective $\ell$ and entails resolving a nested loop optimization, which includes unrolling the entire training trajectory of the relay model optimization (i.e., inner loop) and then updates the condensed graph in the outer loop. A general workflow for this bi-level optimization can be formulated as Algorithm 1 and this process can be prohibitively expensive [72]. To alleviate the complexity of the optimization process, various approximation methods have been proposed as shown in Figure LABEL:figpie (b), including gradient matching, trajectory matching, kernel ridge regression, and distribution matching.
1 Input: Original graph $\mathcal{T}$, relay model $f_{\theta}$, iterations {$T_{1}$, $T_{2}$}.
2 Output: Condensed graph $\mathcal{S}$.
3 Initialise $\mathcal{S}$ and $f_{\theta}$.
[…]
5       Update $\mathcal{S}$ according to $\mathcal{L}^{\mathcal{T}}$.
[…]
7             Optimize $f_{\theta}$ according to $\mathcal{L}^{\mathcal{S}}$.
8
9
10Return: Condensed graph $\mathcal{S}$.
Algorithm 1 The bi-level optimization workflow of GC.
TABLE I: The comparison of the time complexity for diverse optimization objectives. The process in the condensing procedure varies for different GC objectives. Gradient matching includes the gradient calculation. Kernel ridge regression includes the computation of regression solution. Distribution matching includes class representation calculation.
Strategy Pre-processing Condensing procedure
Forward Process Loss Update $\mathcal{S}$ Update $f_{\theta}$
Gradient matching $\mathcal{O}\left(LMd\right)$ $\mathcal{O}\left(\left(N+N^{\prime}\right)dh\right)$ $\mathcal{O}\left(dh\right)$ $\mathcal{O}\left(dh\right)$ $\mathcal{O}\left(N^{\prime}d\right)$ $\mathcal{O}\left(qdh\right)$
Trajectory matching $\mathcal{O}\left(Z(LMd+Ndh)\right)$ $\mathcal{O}\left(qN^{\prime}dh\right)$ N/A $\mathcal{O}\left(dh\right)$ $\mathcal{O}\left(N^{\prime}d\right)$ $\mathcal{O}\left(qdh\right)$
Kernel ridge regression $\mathcal{O}\left(LMd\right)$ $\mathcal{O}\left(\left(N+N^{\prime}\right)dh\right)$ $\mathcal{O}\left(N^{\prime 3}+NCh\right)$ $\mathcal{O}\left(Ch\right)$ $\mathcal{O}\left(N^{\prime}d\right)$ N/A
Distribution matching $\mathcal{O}\left(LMd\right)$ $\mathcal{O}\left(\left(N+N^{\prime}\right)dh\right)$ $\mathcal{O}\left(\left(N+N^{\prime}\right)h\right)$ $\mathcal{O}\left(Ch\right)$ $\mathcal{O}\left(N^{\prime}d\right)$ N/A
Gradient matching. Gradient matching is initially proposed by Zhao et al. [72] and has become the dominant optimization strategy in GC. Specifically, it formulates the condensation objective in Eq. (4) as matching the optimized parameters of the models trained on two datasets:
@@
where $\theta_{0}$ is the initialization of ${\theta}^{\mathcal{T}}$ and ${\theta}^{\mathcal{S}}$. ${\Theta}$ is a specific distribution for relay model initialization. The expectation on $\theta_{0}$ aims to improve the robustness of ${\mathcal{S}}$ to different parameter initialization [39]. $\mathcal{D}\left(\cdot,\cdot\right)$ is the distance measurement. The bi-level objective in Eq. (4) is approximated by matching the model gradients at each training step $t$. In this way, the training trajectory on condensed data can mimic that on the original data, i.e., the models trained on these two datasets converge to similar solutions. The optimization objective $\mathcal{L}_{cond}$ in Eq. (2) is defined as:
$\begin{split}&\mathcal{L}_{cond}=\mathbb{E}_{\theta_{0}\sim{\Theta}}\left[\sum_{t=1}^{T}\mathcal{D}\left(\nabla_{\theta}\mathcal{L}^{\mathcal{T}}\left({\theta_{t}}\right),\nabla_{\theta}\mathcal{L}^{\mathcal{S}}\left({\theta_{t}}\right)\right)\right]\,\,\\ &\text{s.t.}\,\,{\theta}_{t+1}=\operatorname{opt}\left(\mathcal{L}^{\mathcal{S}}\left({\theta_{t}}\right)\right),\end{split}$ (6)
@@
Trajectory matching. Gradient matching primarily aligns with single-step gradients, yet it may accumulate errors when the relay model is iteratively updated using condensed data across multiple steps. To mitigate this problem, Cazenavette et al. [73] propose a multi-step matching approach to approximate Eq. (5) known as trajectory matching.
Trajectory matching trains two separate relay models on condensed data and original data, respectively. These two relay models are trained for distinct steps starting from the same initialization $\theta^{\prime}_{t}$, which is sampled from the original data training checkpoints ${\Theta^{\prime}}$. Then, trajectory matching aims to minimize the discrepancy between the final points of these two training trajectories:
@@
Kernel ridge regression. Gradient matching and trajectory matching approximately optimize the objective Eq. (5) by matching one-step gradient or multi-step trajectory. However, the complex bi-level optimization and inexact solution inevitably lead to slow convergence and performance drop. In light of these challenges, Nguyen et al. [74] transform the original task into a regression problem by substituting the neural network with kernel ridge regression (KRR). This approach enables the direct resolution of the inner loop optimization in Eq. (4) using a closed-form solution. Consequently, the complex nested optimization is simplified into a first-order process by considering the condensed data as the support set and the original data as the target set. The optimization problem described in Eq. (4) is reformulated into a regression task:
@@
Distribution matching. All aforementioned optimization strategies stem from the objective of aligning model performance as outlined in Eq. (4), inevitably introducing the relay model training procedure within the optimization process. To avoid this issue, Zhao et al. [75] propose an alternative optimization principle known as distribution matching, which focuses on matching the feature distributions of condensed and original data. The objective of distribution matching is to generate condensed data with a feature distribution closely approximating that of the original data. Accordingly, the optimization objective of distribution matching is formulated as:
@@
To address the discrepancy in sample quantities between the original and condensed dataset, the distance measurement $\mathcal{D}\left(\cdot,\cdot\right)$ is specifically designed for class-wise comparison, necessitating the use of class labels in distribution matching.
TABLE II: Summary of existing graph condensation research, organized chronologically by publication date. In optimization strategy: “GM” represents gradient matching, “DM” refers to distribution matching, “TM” is short for trajectory matching, “KRR” is kernel ridge regression, and “CTP” denotes computation tree preservation. In relay model: “GNN” is graph neural network, “GNTK” refers to graph neural tangent kernel and “SD” is short for spectral decomposition. In downstream task: “NC” represents node classification, “GC” refers to graph classification, “AD” is anomaly detection and “LP” is short for link prediction. The “Enhanced Module” identifies the specific target for improvements within the GC framework.
Method Category Input Optimization Strategy Output Downstream Task Enhanced Module
Preproc. Objective Model Bi-level Structure construction Sparsification
GCond [25] Effectiveness GM GNN Generative model Threshold NC N/A
DosCond [52] Efficiency GM GNN Parameterization Gumbel softmax GC Optimization
GCDM [63] Efficiency DM GNN Generative model Threshold NC Objective
HCDC [10] Application GM GNN Generative model Threshold NC Condensed graph
MSGC [48] Effectiveness GM GNN Pre-defined structure N/A NC Condensed graph
SGDD [30] Generalization GM GNN Generative model Regularization NC, AD, LP Objective
SFGC [28] Effectiveness TM GNN Pre-defined structure N/A NC Objective
FGD [49] Fairness GM GNN Generative model Threshold NC Objective
GCARe [50] Fairness GM GNN Generative model Threshold NC Relay model
MCond [26] Application GM GNN Generative model Threshold NC Objective
KiDD [29] Efficiency KRR GNTK Parameterization Gumbel softmax GC Relay model
CaT [27] Application DM GNN Pre-defined structure N/A NC Optimization
CTRL [51] Effectiveness GM GNN Generative model Threshold NC, GC Objective
FedGKD [56] Application GM GNN Generative model Gumbel softmax NC Condensed graph
GCEM [64] Generalization DM SD Parameterization N/A NC Relay model
GC-SNTK [61] Efficiency KRR GNTK Pre-defined structure N/A NC Objective
Mirage [69] Efficiency CTP Free Parameterization N/A GC Optimization
GroC [47] Effectiveness GM GNN Generative model Threshold NC, GC Objective
PUMA [65] Application DM GNN Pre-defined structure N/A NC Objective
DisCo [67] Efficiency DM MLP Generative model N/A NC Objective
GEOM [59] Effectiveness TM GNN Generative model Threshold NC Objective
HGCond [53] Application GM GNN Generative model N/A NC Condensed graph
EXGC [54] Efficiency GM GNN Generative model Threshold NC, GC Optimization
OpenGC [62] Generalization KRR GNN Pre-defined structure N/A NC Objective
GCSR [60] Effectiveness TM GNN Parameterization N/A NC Objective
SimGC [68] Efficiency DM GNN Generative model Threshold NC Optimization
FedGC [55] Application GM GNN Generative model Threshold NC Objective
CGC [66] Efficiency DM GNN Parameterization Threshold NC Optimization
RobGC [58] Robustness GM GNN Generative model Threshold NC Condensed graph
TinyGraph [57] Effectiveness GM GNN Generative model Threshold NC Condensed graph
Discussion. The gradient matching approach focuses on aligning short-range model parameters, while trajectory matching enhances data quality by matching long-range model trajectories. Both methods employ the bi-level optimization framework, requiring multiple forward and backward computations and thus incurring high costs and time consumption. In contrast, KRR and distribution matching adopt the single-level optimization framework, leading to more efficient optimization processes. Although distribution matching uses the feature space as a matching proxy and bypasses relay model optimization, this strategy inherently confines it to class-wise comparisons. Consequently, this design restricts its adaptability to diverse task-specific objectives and its applicability to data lacking class labels.
For an intuitive comparison, we present the detailed time complexity of various optimization strategies in Table I. Time complexities for both the pre-processing and condensing phases are assessed separately and the condensing procedure is further divided into forward propagation, process execution, loss calculation, condensed graph updating, and relay model updating. Following GCond [25], the relay model $f_{\theta}$ for all strategies is a $L$ layer SGC [16] incorporating a linear model with the hidden dimension denoted by $h$. For the original graph, $N$, $M$, $d$, and $C$ are the number of nodes, edges, feature dimensions, and classes, respectively. The number of nodes in the condensed graph is represented by $N^{\prime}$. For a fair comparison, the pre-defined adjacency matrix [28] is utilized across all methods. The pre-processing stage for gradient matching [25], kernel ridge regression [62] and distribution matching [46] incorporates non-parametric graph convolution. Trajectory matching [28] entails the pre-training of teacher models and the model quantity is denoted by $Z$. The process execution stage involves different operations specific to each strategy. Specifically, gradient matching entails calculating the gradient w.r.t the relay model parameters twice. Kernel ridge regression calculates the closed-form solution, and distribution matching involves computing the class representation. As for the update of the relay model, gradient matching and trajectory matching necessitate updating the relay model on the condensed graph $q$ times to generate model parameters and trajectories in the inner loop, respectively. Conversely, kernel ridge regression and distribution matching simplify the process by eliminating bi-level optimization, thereby obviating the need to update the relay model.